# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 3 Overview and Categorization
Spatio-Temporal Data [ Unconditional
Diffusion Models
(subsection 4.1 ) [ Probability-based
Diffusion Models
(Sec. 4.1.1) [ Predictive Tasks
(subsection 5.1  subsection 5.4 ) [ Time Series
(subsection 6.1 ) [ Healthcare: DiffEEG [18] DiffECG [74],leaf, text width=22em ] [ Energy: Capel et al. [14],leaf, text width=22em ] [ General: TimeGrad [13] D3VAE [86],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Climate: SwinRDM [16] SEEDS [84],leaf, text width=22em ] [ Traffic: DVGNN [59] SpecSTG [88],leaf, text width=22em ] [ Others: DSTPP [70] DYffusion [29],leaf, text width=22em ] ] ] [ Generative Tasks
(subsection 5.2  subsection 5.3 ) [ Time Series
(subsection 6.1 ) [ Healthcare: Kuo et al. [19] Diff-E [89],leaf, text width=22em ] [ Audio: WaveGrad [20] DiffWave [91],leaf, text width=22em ] [ Recommendation: DiffuASR [23] DiffuRec [95],leaf, text width=22em ] [ Others: TransFusion [24] DiffCharge [96],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Traffic: TrajGDM [60],leaf, text width=22em ] [ General: Yun et al. [69] Wang et al.[25],leaf, text width=22em ] ] ] [ Other Tasks [ Time Series
(subsection 6.1 ) [ Classification: CARD [98],leaf, text width=22em ] [ Anomaly Detection: DDMT [99],leaf, text width=22em ] [ Denoising: Duan et al. [100],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Alignment: ERDiff [25],leaf, text width=22em ] [ Anomaly Detection: DiffTAD [97],leaf, text width=22em ] [ Imputation: Imputation as Inpainting [69],leaf, text width=22em ] ] ] ] [ Score-based
Diffusion Models
(Sec. 4.1.2) [ Predictive Tasks
(subsection 5.1  subsection 5.4 ) [ Time Series
(subsection 6.1 ) [ General: ScoreGrad [71] Biloš et al. [104],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Climate: Hatanaka et al. [27],leaf, text width=22em ] [ General: DYffusion [29],leaf, text width=22em ] ] ] [ Generative Tasks
(subsection 5.2  subsection 5.3 ) [ Time Series
(subsection 6.1 ) [ Audio: SGMSE [32] StoRM [105],leaf, text width=22em ] [ Healthcare: DeScoD-ECG [57],leaf, text width=22em ] [ Others: SaSDim [72] Crabbé et al. [106],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Healthcare: EHRDiff [30],leaf, text width=22em ] ] ] ] ] [ Conditional
Diffusion Models
(Sec. 4.2.1) [ Probability-based
Diffusion Models
(Sec. 4.2.1) [ Predictive Tasks
(subsection 5.1  subsection 5.4 ) [ Time Series
(subsection 6.1 ) [ AIOps: Maat [37],leaf, text width=22em ] [ Manufacture: Zhang et al. [108] SCDDPM [181],leaf, text width=22em ] [ Energy: VGCDM [263] Dong at al. [107],leaf, text width=22em ] [ Recommendation: Yang et al. [167],leaf, text width=22em ] [ Others: TimeDiff [34] Wang et al.  [262],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Traffic: DiffUFlow [35] DiffSTG [43],leaf, text width=22em ] [ Climate: DiffESM [15] DiffSTG [43],leaf, text width=22em ] [ General: Hu et al. [170],leaf, text width=22em ] ] ] [ Generative Tasks
(subsection 5.2  subsection 5.3 ) [ Time Series
(subsection 6.1 ) [ Healthcare: BioDiffusion [51] MEDiC [110], leaf, text width=22em ] [ Finance: DiffsFormer [52] ,leaf, text width=22em ] [ Audio: CDiffuSE [175] DOSE [177],leaf, text width=22em ] [ Others: Yan et al. [53] DiffTime [117],leaf, text width=22em ] [ General: Time Weaver [182] Shirzad at al. [183],leaf, text width=22em ] ] [ Spatio-Temporal Data
(subsection 6.2 ) [ Traffic: DiffTraj [81] ControlTraj [290],leaf, text width=22em ] [ Others: PriSTI [111],leaf, text width=22em ] ] ] ] [ Score-based
Diffusion Models
(Sec. 4.2.1) [ Generative Tasks
(subsection 5.2  subsection 5.3 ) [ Time Series
(subsection 6.1 ) [ Healthcare: Seki et al. [54],leaf, text width=22em ] [ Audio: UNIVERSE [180],leaf, text width=22em ] [ General: CSDI [56] Lim et al.[116],leaf, text width=22em ] ] ] ] ] ]
@@
Figure 7: A comprehensive taxonomy of diffusion models for time series and spatio-temporal data, categorized according to methodologies (i.e., unconditional vs. conditional), tasks (e.g., predictive versus generative), data types, and applications.
@@
This section presents an overview and classification of diffusion models for addressing challenges in time series and spatio-temporal data analysis. Our survey organizes the discussion along four primary dimensions: categories of diffusion models, types of tasks, data modalities, and practical applications. A comprehensive summary of notable related works is depicted in Figure 7 . We categorize the existing literature into two primary groups: unconditioned and conditioned diffusion models, focusing on time series and spatio-temporal data.
@@
In the unconditioned category, diffusion models operate in an unsupervised manner to generate data samples without the need for supervision signals. This setting represents the foundational approach for analyzing time series and spatio-temporal data. Within this category, literature can be further divided into probability-based and score-based diffusion models. Examples include denoised diffusion probabilistic models (DDPMs) [2] and score-based stochastic differential equations (Score SDEs) [6, 4], as introduced in section 2 . Research in this category is broadly organized into two task groups: predictive and generative tasks. Predictive tasks typically involve forecasting and anomaly detection, leveraging historical data and patterns to anticipate current and/or future events. Generative tasks, conversely, focus on identifying patterns within extensive datasets to generate new content, such as time series imputation and augmentation. Methods are developed for both primary data modalities: time series and spatio-temporal data, catering to a wide range of applications across various sectors, including healthcare, energy, climate, traffic, and more.
@@
In the conditioned category, diffusion models are tailored for conditioned analysis of time series and spatio-temporal data. Empirical studies have shown that conditional generative models, which utilize data labels, are easier to train and yield superior performance compared to their unconditional counterparts [75]. In this context, labels (a.k.a. conditions) often derive from various sources, such as extracted short-term trends [34] and urban flow maps [35], to enhance model inferences. This category embraces both probability-based and score-based diffusion models for predictive and generative tasks, offering a fresher perspective on leveraging diffusion models to tackle practical challenges in time series and spatio-temporal data analysis under specific constraints.
@@
Building on the foundational understanding of model categories, task types, data modalities, and application domains, we delve deeper into the exploration of diffusion models for time series and spatio-temporal data analysis across the following sections. Each section is designed to unpack the complexities and nuances inherent in the application of diffusion models, providing a comprehensive overview from multiple perspectives. In section 4 , we explore diffusion model landscapes, highlighting distinctions between unconditioned and conditioned approaches and their implications. section 5 analyzes tasks from predictive and generative viewpoints, detailing specific functions such as forecasting, generation, anomaly detection, and data imputation. section 6 examines data modalities, differentiating between time series and spatio-temporal data to outline model challenges and applicability. Lastly, section 7 extends the discussion to application fields, demonstrating diffusion models’ utility across sectors like healthcare, traffic, sequential recommendation, climate, energy, and audio. This structured exploration aims to equip readers with an in-depth understanding of the potential and current state of diffusion models for addressing complex time series and spatio-temporal data challenges.
4 Model Perspective
@@
In this section, we will analyze how to use diffusion models for time series and spatio-temporal data from a model perspective. Specifically, we will focus on standard diffusion models (DDPM and score SDE) and improved diffusion models (conditional diffusion model, LDM, DDIM, and others).
4.1 Standard Diffusion Model
@@
Standard diffusion models include probability-based model, i.e., DDPM and score-based model, score SDE, which have been described in detail in Sec. 2.3.1 and Sec. 2.3.2, respectively. Currently, diffusion models based on these two are also the most common methods for time series and spatio-temporal data analysis.
4.1.1 Probability-Based Model
@@
Based on the probabilistic-based standard model, we mainly introduce DDPM there, which uses a discrete framework, decomposing the diffusion process into a fixed number of steps. This method adds noise to the data at discrete intervals and then learns to reverse this process, generating data from noise. Both the training and inference process of the model are conducted on these discrete steps, making the process essentially stepwise and iterative.
@@
TimeGrad is among the most classic DDPM-based time series forecasting methods, identified as a score-matching model that, through rigorous validation, has been proven effective in real-world datasets comprising thousands of interconnected dimensions [13]. D3VAE proposes a bidirectional variational auto-encoder that integrates diffusion, denoising, and disentanglement to enhance time series data without introducing extraneous uncertainties. This model marries multi-scale denoising score matching with bi-directional variational autoencoding for forecasting tasks [86]. In a similar vein, TSDiff leverages implicit probability densities to iteratively refine the base forecasts, thereby addressing three distinct tasks: forecasting, refining, and generating synthetic data [164]. For the universal challenge of anomaly detection, Pintilie et al. have proposed two multivariate time series anomaly detection algorithms based on diffusion models, securing leading outcomes [259]. Concurrently, D3R tackles temporal anomaly detection drifts through decomposition and reconstruction, employing data-time mix-attention for dynamic decomposition alongside diffusion models for end-to-end training, thus overcoming the limitations imposed by local sliding windows and unstable data scenarios [261]. CARD presents a denoising diffusion-based generative model alongside a pretrained conditional mean estimator to serve both classification and regression tasks [98]. TS-Diffusion is specifically designed for complex sequences marked by irregular sampling, missingness, and extensive temporal feature dimensions, introducing a comprehensive model consisting of an ODE encoder, a representational learning module, and an ODE decoder to handle such intricate time series [278]. Furthermore, WaveGrad [20], DiffWave [91], and DiffuASR [23] have each successfully applied DDPM to waveforms, audio generation, and sequence recommendation, respectively.
@@
On the other hand, in the field of spatio-temporal data, significant advances have been made with DDPM predominantly starting from 2023, indicating an exciting opportunity for more applications and theoretical breakthroughs. Yun et al. have put forward ”Imputation as Inpainting,” which integrates an unconditional diffusion model based on graph neural networks to first forecast complete spatio-temporal data. This approach adjusts the generation process by sampling in unobserved areas using the information from observed data, thus estimating missing values in spatio-temporal data [69]. For traffic and specifically trajectory forecasting, the adoption of graph-based DDPM methods is becoming more common. For example, MID++ leverages a GNN-based ALEncoder to extract spatio-temporal features and improves the training process with an important sampling strategy, leading to better trajectory forecasting [279]. In a similar vein, DiffTAD incorporates a Transformer-based decoupled time and space encoder to model spatial interactions among vehicles and conducts anomaly detection by evaluating the differences between query trajectories and their reconstructions [83]. Moreover, SpecSTG tackles the challenges of insufficient representation of spatial network features and the inability to detect unexpected fluctuations in future observations in traffic flow forecasting. It achieves this by converting the learning process to the spectral domain, generating Fourier representations of future time series that carry spatial information [88].
4.1.2 Score-Based Model
@@
ScoreGrad and TimeGrad share similar goals and can be described as SDE-based TimeGrad models. They expand the diffusion process into a continuous spectrum and employ time series feature extraction modules along with a condition-based stochastic differential equation for score matching to facilitate forecasting [71]. This structure finds parallels in the work of Biloš et al. [104]. Additionally, recent studies by Crabbé et al. delve into representing time series data in the frequency domain using SDEs, showcasing how denoising score matching methods can facilitate diffusion modeling within the frequency domain based on differing time-frequency spaces [106]. TimeADDM introduces a score-based diffusion model for unsupervised anomaly detection in multivariate time series, applying diffusion steps to representations that encapsulate global time correlations through recurrent embeddings and designing a suite of reconstruction strategies to compute anomaly scores at various diffusion intensities [282]. Moreover, Li et al. extensively discuss the impact of noise samples in SDEs on data quality and the corresponding robustness of models [280].
@@
From an application standpoint, score-based diffusion models have shown great promise in continuous data fields such as healthcare and acoustic data. EHRDiff [30] and DeScoD-ECG [57] leverage score-based diffusion models for realistic EHR synthesis and ECG data generation, respectively, incorporating algorithms or structures pertinent to the features of medical data, thus improving applicability in real-world settings. In audio-related applications, the focus has been on speech enhancement and dereverberation, with initial proposals by Welker et al. [284] and Richter et al. [32] for employing score-based diffusion models for speech enhancement, transforming time domain speech data into time-frequency information, and treating them as image input for models. StoRM introduces a stochastic resampling method, utilizing forecasting from a predictive model as guidance for subsequent diffusion steps, achieving higher-quality samples under complex conditions with high signal-to-noise ratios using fewer diffusion steps and more streamlined models [105]. Similarly, Lay et al. emphasize improving model efficiency and optimizing model parameters, proposing a Brownian bridge-based forward process to reconcile the gap between the forward process’s end distribution and the prior distribution used in inference for the reverse process [178].
@@
In the realm of spatio-temporal data, methods remain scarce, with Sasdim and DYffusion standing out. Sasdim, an adaptive noise scaling diffusion model, effectively performs spatial-temporal data imputation by capturing dynamic spatio-temporal dependencies through novel loss functions and global spatio-temporal convolution modules [72]. DYffusion, a dynamics-informed diffusion model for spatio-temporal data forecasting, distinguishes itself from probabilistic models by integrating temporal dynamics directly into the diffusion steps and training a stochastic, time-conditioned interpolator alongside a predictor network to simulate the forward and reverse processes of the standard diffusion model, thus striking a balance between performance and efficiency [285].
4.2 Improved/Advanced Diffusion Model

### v5

@@ 3. Overview and Categorization
Spatio-Temporal Data
Unconditional Diffusion Models (subsection 4.1 ) Probability-based Diffusion Models (Sec. 4.1.1) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) Healthcare: DiffEEG (Shu et al., 2023) DiffECG (Neifar et al., 2023) General: TimeGrad (Rasul et al., 2021) D3VAE (Li et al., 2022a) Spatio-Temporal Data (subsection 6.2 ) Climate: SwinRDM (Chen et al., 2023a) SEEDS (Li et al., 2023a) Traffic: DVGNN (Liang et al., 2023) SpecSTG (Lin et al., 2024) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: Kuo et al. (Nicholas et al., 2023) Diff-E (Kim et al., 2023a) Audio: WaveGrad (Chen et al., 2020) DiffWave (Kong et al., 2020) Recommendation: DiffuASR (Liu et al., 2023g) DiffuRec (Li et al., 2023d) Spatio-Temporal Data (subsection 6.2 ) Traffic: TrajGDM (Chu et al., 2024) General: Yun et al. (Yun et al., 2023) Wang et al.(Wang et al., 2024b) Other Tasks Time Series (subsection 6.1 ) Classification: CARD (Han et al., 2022) Anomaly Detection: DDMT (Yang et al., 2023b) Spatio-Temporal Data (subsection 6.2 ) Anomaly Detection: DiffTAD (Li et al., 2024b) Imputation: Imputation as Inpainting (Yun et al., 2023) Score-based Diffusion Models (Sec. 4.1.2) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) General: ScoreGrad (Yan et al., 2021) Biloš et al. (Biloš et al., 2022) Spatio-Temporal Data (subsection 6.2 ) Climate: Hatanaka et al. (Hatanaka et al., 2023) General: DYffusion (Cachay et al., 2023) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Audio: SGMSE (Richter et al., 2023) StoRM (Lemercier et al., 2023) Healthcare: DeScoD-ECG (Li et al., 2023b) Spatio-Temporal Data (subsection 6.2 ) Healthcare: EHRDiff (Yuan et al., 2023c) Conditional Diffusion Models (Sec. 4.2.1) Probability-based Diffusion Models (Sec. 4.2.1) Predictive Tasks (subsection 5.1 subsection 5.4 ) Time Series (subsection 6.1 ) AIOps: Maat (Lee et al., 2023) Manufacture: Zhang et al. (Zhang et al., 2023e) SCDDPM (Chen et al., 2024a) Energy: VGCDM (Liu et al., 2023a) Dong at al. (Dong et al., 2023) Spatio-Temporal Data (subsection 6.2 ) Traffic: DiffUFlow (Zheng et al., 2023) DiffSTG (Wen et al., 2023) Climate: DiffESM (Bassetti et al., 2023) DiffSTG (Wen et al., 2023) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: BioDiffusion (Li et al., 2024d) MEDiC (Sharma et al., 2023) Finance: DiffsFormer (Gao et al., 2024) Audio: CDiffuSE (Lu et al., 2022a) DOSE (Tai et al., 2024) Spatio-Temporal Data (subsection 6.2 ) Traffic: DiffTraj (Zhu et al., 2024a) ControlTraj (Zhu et al., 2024b) Others: PriSTI (Liu et al., 2023c) Score-based Diffusion Models (Sec. 4.2.1) Generative Tasks (subsection 5.2 subsection 5.3 ) Time Series (subsection 6.1 ) Healthcare: Seki et al. (Seki et al., 2023) Audio: UNIVERSE (Serrà et al., 2022)
@@
Figure 6. A comprehensive taxonomy of diffusion models for time series and spatio-temporal data, categorized according to methodologies (i.e., unconditional vs. conditional), tasks (e.g., predictive versus generative), data types, and applications.
@@
This section presents an overview and classification of diffusion models for addressing challenges in time series and spatio-temporal data analysis. Our survey organizes the discussion along four primary dimensions: categories of diffusion models, types of tasks, data modalities, and practical applications [^2] All summarized and categorized papers document: https://docs.google.com/spreadsheets/d/1X5ujA-yXCzkzz4Uu9ErP8mu_Jg2zWk4fwZrwrMcqMoI.. A comprehensive summary of notable related works is depicted in Figure 6 . We categorize the existing literature into two primary groups: unconditioned and conditioned diffusion models, focusing on time series and spatio-temporal data.
@@
In the unconditioned category, diffusion models operate in an unsupervised manner to generate data samples without the need for supervision signals. This setting represents the foundational approach for analyzing time series and spatio-temporal data. Within this category, literature can be further divided into probability-based and score-based diffusion models. Examples include denoised diffusion probabilistic models (DDPMs) (Ho et al., 2020) and score-based stochastic differential equations (Score SDEs) (Song et al., 2020b; Song and Ermon, 2019), as introduced in section 2 . Research in this category is broadly organized into two task groups: predictive and generative tasks. Predictive tasks typically involve forecasting and anomaly detection, leveraging historical data and patterns to anticipate current and/or future events. Generative tasks, conversely, focus on identifying patterns within extensive datasets to generate new content, such as time series imputation and augmentation. Methods are developed for both primary data modalities: time series and spatio-temporal data, catering to a wide range of applications across various sectors, including healthcare, energy, climate, traffic, and more.
@@
In the conditioned category, diffusion models are tailored for conditioned analysis of time series and spatio-temporal data. Empirical studies have shown that conditional generative models, which utilize data labels, are easier to train and yield superior performance compared to their unconditional counterparts (Bao et al., 2022). In this context, labels (a.k.a. conditions) often derive from various sources, such as extracted short-term trends (Shen and Kwok, 2023) and urban flow maps (Zheng et al., 2023), to enhance model inferences. This category embraces both probability-based and score-based diffusion models for predictive and generative tasks, offering a fresher perspective on leveraging diffusion models to tackle practical challenges in time series and spatio-temporal data analysis under specific constraints.
@@
Building on the foundational understanding of model categories, task types, data modalities, and application domains, we delve deeper into the exploration of diffusion models for time series and spatio-temporal data analysis across the following sections. Each section is designed to unpack the complexities and nuances inherent in the application of diffusion models, providing a comprehensive overview from multiple perspectives. In section 4 , we explore diffusion model landscapes, highlighting distinctions between unconditioned and conditioned approaches and their implications. section 5 analyzes tasks from the predictive and generative viewpoints, detailing specific functions such as forecasting, generation, anomaly detection, and data imputation. section 6 examines the data modalities, differentiating between time series and spatio-temporal data to outline model challenges and applicability. Lastly, section 7 extends the discussion to application fields, demonstrating diffusion models’ utility across sectors like healthcare, traffic, sequential recommendation, climate, energy, and audio. This structured exploration aims to equip readers with an in-depth understanding of the potential and current state of diffusion models for addressing complex time series and spatio-temporal data challenges.
4. Model Perspective
@@
In this section, we will analyze how to use diffusion models for time series and spatio-temporal data from a model perspective. Specifically, we will focus on standard diffusion models (DDPM and score SDE) and improved diffusion models (conditional diffusion model, LDM, DDIM, and others). In addition to the above content, we will also discuss in detail an important issue in the model section, namely, the denoisers in the diffusion model.
4.1. Standard Diffusion Model
@@
Standard diffusion models include probability-based model, i.e., DDPM and score-based models, score SDE, which have been described in detail in Sec. 2.3.1 and Sec. 2.3.2, respectively. Currently, diffusion models based on these two are also the most common methods for time series and spatio-temporal data analysis.
4.1.1. Probability-Based Model
@@
Based on the probabilistic-based standard model, we mainly introduce DDPM there, which uses a discrete framework, decomposing the diffusion process into a fixed number of steps. This method adds noise to the data at discrete intervals and then learns to reverse this process, generating data from noise. Both the training and inference process of the model are conducted in these discrete steps, making the process essentially stepwise and iterative.
@@
TimeGrad is among the most classic DDPM-based time series forecasting methods, identified as a score-matching model that, through rigorous validation, has been proven effective in real-world datasets comprising thousands of interconnected dimensions (Rasul et al., 2021). D3VAE proposes a bidirectional variational auto-encoder that integrates diffusion, denoising, and disentanglement to enhance time series data without introducing extraneous uncertainties. This model marries multi-scale denoising score matching with bi-directional variational autoencoding for forecasting tasks (Li et al., 2022a). In a similar vein, TSDiff leverages implicit probability densities to iteratively refine the base forecasts, thereby addressing three distinct tasks: forecasting, refining, and generating synthetic data (Kollovieh et al., 2024). For the universal challenge of anomaly detection, Pintilie et al. have proposed two multivariate time series anomaly detection algorithms based on diffusion models, securing leading outcomes (Pintilie et al., 2023). Concurrently, D3R tackles temporal anomaly detection drifts through decomposition and reconstruction, employing data-time mix-attention for dynamic decomposition alongside diffusion models for end-to-end training, thus overcoming the limitations imposed by local sliding windows and unstable data scenarios (Wang et al., 2024c). CARD presents a denoising diffusion-based generative model alongside a pretrained conditional mean estimator to serve both classification and regression tasks (Han et al., 2022). TS-Diffusion is specifically designed for complex sequences marked by irregular sampling, missingness, and extensive temporal feature dimensions, introducing a comprehensive model consisting of an ODE encoder, a representational learning module, and an ODE decoder to handle such intricate time series (Li, 2023). Furthermore, WaveGrad (Chen et al., 2020), DiffWave (Kong et al., 2020), and DiffuASR (Liu et al., 2023g) have each successfully applied DDPM to waveforms, audio generation, and sequence recommendation, respectively.
@@
On the other hand, in the field of spatio-temporal data, significant advances have been made with DDPM predominantly starting from 2023, indicating an exciting opportunity for more applications and theoretical breakthroughs. Yun et al. have put forward ”Imputation as Inpainting,” which integrates an unconditional diffusion model based on graph neural networks to first forecast complete spatio-temporal data. This approach adjusts the generation process by sampling in unobserved areas using the information from observed data, thus estimating missing values in spatio-temporal data (Yun et al., 2023). For traffic and specifically trajectory forecasting, the adoption of graph-based DDPM methods is becoming more common. For example, MID++ leverages a GNN-based ALEncoder to extract spatio-temporal features and improves the training process with an important sampling strategy, leading to better trajectory forecasting (Yao et al., 2023b). In a similar vein, DiffTAD incorporates a Transformer-based decoupled time and space encoder to model spatial interactions among vehicles and conducts anomaly detection by evaluating the differences between query trajectories and their reconstructions (Li et al., 2024a). Moreover, SpecSTG tackles the challenges of insufficient representation of spatial network features and the inability to detect unexpected fluctuations in future observations in traffic flow forecasting. It achieves this by converting the learning process to the spectral domain, generating Fourier representations of future time series that carry spatial information (Lin et al., 2024).
4.1.2. Score-Based Model
@@
ScoreGrad and TimeGrad share similar goals and can be described as SDE-based TimeGrad models. They expand the diffusion process into a continuous spectrum and employ time series feature extraction modules along with a condition-based stochastic differential equation for score matching to facilitate forecasting (Yan et al., 2021). This structure finds parallels in the work of Biloš et al. (Biloš et al., 2022). Additionally, recent studies by Crabbé et al. delve into representing time series data in the frequency domain using SDEs, showcasing how denoising score matching methods can facilitate diffusion modeling within the frequency domain based on differing time-frequency spaces (Crabbé et al., 2024). TimeADDM introduces a score-based diffusion model for unsupervised anomaly detection in multivariate time series, applying diffusion steps to representations that encapsulate global time correlations through recurrent embeddings and designing a suite of reconstruction strategies to compute anomaly scores at various diffusion intensities (Hu et al., 2024). Moreover, Li et al. extensively discuss the impact of noise samples in SDEs on data quality and the corresponding robustness of the models (Li et al., 2024c).
@@
From an application standpoint, score-based diffusion models have shown great promise in continuous data fields such as healthcare and acoustic data. EHRDiff (Yuan et al., 2023c) and DeScoD-ECG (Li et al., 2023b) leverage score-based diffusion models for realistic EHR synthesis and ECG data generation, respectively, incorporating algorithms or structures pertinent to the features of medical data, thus improving applicability in real-world settings. In audio-related applications, the focus has been on speech enhancement and dereverberation, with initial proposals by Welker et al. (Welker et al., 2022) and Richter et al. (Richter et al., 2023) for employing score-based diffusion models for speech enhancement, transforming time domain speech data into time-frequency information, and treating them as image input for models. StoRM introduces a stochastic resampling method, utilizing forecasting from a predictive model as guidance for subsequent diffusion steps, achieving higher-quality samples under complex conditions with high signal-to-noise ratios using fewer diffusion steps and more streamlined models (Lemercier et al., 2023). Similarly, Lay et al. emphasize improving model efficiency and optimizing model parameters, proposing a Brownian bridge-based forward process to reconcile the gap between the forward process’s end distribution and the prior distribution used in inference for the reverse process (Lay et al., 2023).
@@
In the realm of spatio-temporal data, methods remain scarce, with Sasdim and DYffusion standing out. Sasdim, an adaptive noise scaling diffusion model, effectively performs spatio-temporal data imputation by capturing dynamic spatio-temporal dependencies through novel loss functions and global spatio-temporal convolution modules (Zhang et al., 2023d). DYffusion, a dynamics-informed diffusion model for spatio-temporal data forecasting, distinguishes itself from probabilistic models by integrating temporal dynamics directly into the diffusion steps and training a stochastic, time-conditioned interpolator alongside a predictor network to simulate the forward and reverse processes of the standard diffusion model, thus striking a balance between performance and efficiency (Rühling Cachay et al., 2024).
4.2. Improved/Advanced Diffusion Model